OpenStudy (anonymous):

Solve the system of equations: 3x+4y+z=7 2y+z=3 -5x+3y+8z=-31 Can't figure out triangular-system method. Help?

OpenStudy (blank ):

x=-58/49 z=-223/49 y=185/49 Those are the solutions. Tell me if you need the steps. (it is going to be a lot)

OpenStudy (anonymous):

um, please tell me the basic steps to figure it out, if u dont mind, thanks.

OpenStudy (anonymous):

you could prolly upload a pic of your scratch paper with your steps.

OpenStudy (blank ):

3x+4y+z=7 2y+z=3 -5x+3y+8z=-31 Since 2y does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 2y from both sides. z=-2y+3 Since the equation was solved for z, replace all occurrences of z in the other equations with the solution (-2y+3). 3x+4y+(-2y+3)=7 z=-2y+3 -5x+3y+8(-2y+3)=-31 Since 4y and -2y are like terms, add -2y to 4y to get 2y. 3x+2y+3=7 Multiply 8 by each term inside the parentheses. -5x+3y+(-16y+24)=-31 Since 3y and -16y are like terms, add -16y to 3y to get -13y. -5x-13y+24=-31 Move all terms not containing x to the right-hand side of the equation. 3x=-2y+4 Divide each term in the equation by 3. x=-(2(y-2))/(3) Since the equation was solved for x, replace all occurrences of x in the other equations with the solution (-(2(y-2))/(3)). x=-(2(y-2))/(3) z=-2y+3 -5(-(2(y-2))/(3))-13y+24=-31 Multiply -2 by each term inside the parentheses. x=(-2y+4)/(3) Divide each term in the numerator by the denominator. x=-(2y)/(3)+(4)/(3) Multiply -2 by each term inside the parentheses. -5((-2y+4)/(3))-13y+24=-31 Divide each term in the numerator by the denominator. -5(-(2y)/(3)+(4)/(3))-13y+24=-31 Combine all similar expressions. -5((-2y+4)/(3))-13y+24=-31 Multiply -5 by each term inside the parentheses. -(10(y-2))/(3)-13y+24=-31 Multiply -10 by each term inside the parentheses. (-10y+20)/(3)-13y+24=-31 Divide each term in the numerator by the denominator. -(10y)/(3)+(20)/(3)-13y+24=-31 Combine 24+(20)/(3) into a single expression by finding the least common denominator (LCD). The LCD of 24+(20)/(3) is 3. -(10y)/(3)-13y+(92)/(3)=-31 Combine all similar expressions. (-10y+92)/(3)-13y=-31 Factor out the GCF of -2 from -10y+92. (-2(5y-46))/(3)-13y=-31 Move the -1 to the front of the fraction. -(2(5y-46))/(3)-13y=-31 Multiply each term by a factor of 1 that will equate all the denominators. In this case, all terms need a denominator of 3. -(2(5y-46))/(3)-13y*(3)/(3)=-31 Multiply 13y by 3 to get 39y. -(2(5y-46))/(3)-(39y)/(3)=-31 The numerators of expressions that have equal denominators can be combined. In this case, -(2(5y-46))/(3) and -((39y))/(3) have the same denominator of 3, so the numerators can be combined. (-2(5y-46)-(39y))/(3)=-31 Combine all similar expressions in the polynomial. (-49y+92)/(3)=-31 Multiply each term in the equation by 3. -49y+92=-93 Move all terms not containing y to the right-hand side of the equation. -49y=-185 Since the equation was solved for y, replace all occurrences of y in the other equations with the solution ((185)/(49)). y=(185)/(49) Since the equation was solved for y, replace all occurrences of y in the other equations with the solution ((185)/(49)). x=-(2((185)/(49)))/(3)+(4)/(3) z=-2((185)/(49))+3 y=(185)/(49) Remove the single term factors from the expression. x=-(370)/(147)+(4)/(3) Move the minus sign from the numerator to the front of the expression. x=-(174)/(147) Combine (4)/(3)-(370)/(147) into a single expression by finding the least common denominator (LCD). The LCD of (4)/(3)-(370)/(147) is 147. x=-(58)/(49) Multiply -2 by each term inside the parentheses. z=-(370)/(49)+3 Combine 3-(370)/(49) into a single expression by finding the least common denominator (LCD). The LCD of 3-(370)/(49) is 49. z=-(223)/(49) Since the equation was solved for x, replace all occurrences of x in the other equations with the solution (-(58)/(49)). x=-58/49 z=-223/49 y=185/49 Hope this helps. :)

OpenStudy (anonymous):

thank you so much:)

OpenStudy (blank ):

ǝɯoɔlǝʍ ǝɹ,noʎ :)